OpenStudy (anonymous):

Consider the curve x^3 + 3 xy + y^3 = 5 The equation of the tangent line to the curve at the point (1,1) has the form y = mx + b where m = and b =

OpenStudy (freckles):

Well we first need to differentiate both sides of the equation in your first line.

OpenStudy (anonymous):

We do that to so we can find m right? Because the slope would be the derivative

OpenStudy (freckles):

Yes.

OpenStudy (anonymous):

The 5 goes away obviously then, let me see if I can figure out how to do the other side!

OpenStudy (anonymous):

Can you help me differentiate the x^3 + 3 xy + y^3 without giving it away completely to me? :P

OpenStudy (freckles):

Well I think you can differentiate x^3 w.r.t. x pretty easily using power rule

OpenStudy (freckles):

3(xy) 3 is a constant multiple so use constant multiple rule 3(xy) xy is a product so use product rule

OpenStudy (anonymous):

Okay, so just because it's 3 sums added together doesn't effect the sum rule? I've only seen it used with 2 that's why I got confused

OpenStudy (freckles):

\[\frac{d}{dx}(xy)=y \frac{d}{dx}(x)+x \frac{d}{dx}(y) =y \frac{dx}{dx}+x \frac{dy}{dx}=y(1)+x \frac{dy}{dx} \\ \frac{d(xy)}{dx}=y+x \frac{dy}{dx}=y+xy' \]

OpenStudy (freckles):

\[C \cdot \frac{d(xy)}{dx}=C(y+xy')\] where C is a constant

OpenStudy (anonymous):

Let me work on this I think I get it! Thanks for your help!

OpenStudy (anonymous):

Is it wrong to use 3x when doing the product rule instead of using the constant rule? Sorry for all the questions I'm just trying to figure out this stuff completely as I have a midterm tomorrow hahaha

OpenStudy (freckles):

But you still need to use product for (3x)*y

OpenStudy (anonymous):

Yes I got 3x+3y after doing the product rule

OpenStudy (freckles):

:(

OpenStudy (freckles):

derivative of y is y'

OpenStudy (freckles):

(3x*y)'=(3x)'y+(3x)(y)' =3y+(3x)y' =3y+3xy'

OpenStudy (anonymous):

Cant use power rule on that y'? :O

OpenStudy (freckles):

chain rule

OpenStudy (freckles):

y is a function of x so that is why we say derivative of y w.r.t x is y'

OpenStudy (anonymous):

Ohhhhh okay!!!!! Yes

OpenStudy (anonymous):

Thanks a ton I got it from here!